# HFGCN:Hypergraph Fusion Graph Convolutional Networks for Skeleton-Based Action Recognition

### v1

@@ 3.5 Temporal Conv Module
To model the temporal correlations of human pose, we incorporate the Multi-Scale Temporal Convolution (MS-TC) module Liu et al. (2020); Chen et al. (2021) into our final model. The architecture of this module is illustrated in Fig 1(b). It consists of three convolution branches, each starting with a $1\times 1$ convolution to reduce the channel dimension, followed by varying combinations of kernel sizes and dilations. The outputs from these convolution branches are concatenated to form the final temporal feature representation.
@@ 4.1.2 NTU RGB+D 120
Liu et al. (2019) is an extended version of NTU RGB+D, containing an additional 57,367 skeleton sequences across 60 new action classes. Similar to the original dataset, two evaluation benchmarks are recommended: (1) Cross-Subject (X-Sub), where the 106 subjects are divided into training and testing groups, and (2) Cross-Setup (X-Set), where data from samples with even setup IDs are used for training, and data from samples with odd setup IDs are used for testing.
@@ 4.3 Comparison With the State-of-the-Art Methods
In this subsection, we compare our proposed method with a comprehensive list of state-of-the-art skeleton-based action recognition methods across all the three datasets. We adopt a multi-stream fusion strategy, utilizing four different modalities: joint, bone, joint motion, and bone motion. These modalities are explained as follows: The joint modality refers to the original skeleton coordinates, the bone modality represents the differential of spatial coordinates, and the joint motion and bone motion modalities capture the differential along the temporal dimension of the joint and bone modalities, respectively. The results from these four modalities are fused to produce the final outcome.
@@
For NW-UCLA, our proposed method achieved an accuracy of 96.1%, slightly worse than InfoGCN’s 96.6%. This dataset presents particular challenges, as it contains fewer training samples, which makes it harder for our model to perform optimally. Furthermore, the dataset only includes 20 skeleton points, which restricts the effective utilization of hypergraph topology.
Table 2 shows the accuracy of our model across four different modalities and four test benchmarks from two datasets. The results clearly show that using skeleton information is better than using either joint point or motion information alone. It is evident from the table that using skeleton information outperforms using joint point information alone, and incorporating motion information enhances performance across all scenarios. This is because skeleton information encompasses more hierarchical features and is more suitable for directed topological modeling. The fusion of multiple modalities leads to the best performance, further validating the superiority of our method.
@@ 4.4 Ablation Studies
In this subsection, we investigate the impact of different modules in our method on its overall performance. We use ST-GCNYan et al. (2018) as the baseline and analyze the contributions of individual modules by sequentially adding them. For the ablation study, all experiments are conducted on the X-Sub benchmark of NTU RGB+D 120, using only the joint modality as input.
@@ 4.4.1 Hypergraph Topology Studies
For the hypergraph topology we designed, we investigated whether different topological structures might yield varying different accuracy rates for identifying different actions. We conduct experiments on three distinct hypergraph topologies and their pairwise combinations. In addition, we compare the model parameter size and recognition accuracy across different configurations, including fusions of data from various modalities. The experimental results are presented in Table 3.
As shown in the table, the model that integrates all three topological structures achieves the best recognition performance.Among the individual topologies, h3 demonstrates the highest accuracy when all four modalities are fused. For the two-topology combinations, the pairing of h1 and h3 achieve the best accuracy. Nonetheless, the highest recognition accuracy is consistently obtained when all three hypergraph topologies are combined.
@@ 4.4.2 Hypergraph Module Studies
To verify the effectiveness of the hypergraph module and the contribution of different components in our method, we compare our hypergraph module with configurations that exclude the hypergraph component. And we analyze the size and accuracy of our network with and without the hypergraph module. The comparison includes four configurations: the GC module with hypergraph (HGC) and without (GC), as well as the Attention module with hypergraph (HA) and without (A).
As shown in the Table 4, the network’s recognition performance improves progressively with the addition of components. The results demonstrate that incorporating the hypergraph module in our method significantly enhances the baseline model’s recognition performance. Specifically, the inclusion of our proposed HGCM and HAM modules improves accuracy by approximately 1.6%. Although the hypergraph structure increases the model’s parameter count, the overall model size remains compact, and the accuracy gains are substantial.
Notably, our hypergraph attention and hypergraph convolution modules capture the high-order relationships between different skeletal points, highlighting the advantages of using high-order relationships for over point pairs in topological modeling. These results validate the effectiveness of our method, which leverages hypergraph topology for high-order modeling between bone points, leading to substantial improvements in performance.
@@ 5 Conclusion
In this paper, we propose a Hypergraph-Fused Graph Convolutional Network (HFGCN) for action recognition. We design a hypergraph topology structure tailored to the characteristics of human motion and employ a Hypergraph Attention Module (HAM) to learn the high-order relationships between skeleton points. This enables the model to focus on the correlations within skeleton point groups and optimize the topological modeling at the frame level. Furthermore, we introduce a Hypergraph Convolution Module (HGCM) to refine the correlation between skeleton points and skeleton point groups at the channel level. Experimental results on the NTU RGB+D and NTU RGB+D 120 datasets demonstrate that the proposed HFGCN outperforms existing skeleton-based methods. However, the hypergraph in our model is manually designed based on human motion theory, and future work should focus on optimizing this aspect for further improvement.

### v3

@@ 3.5 Temporal Conv Module
To model the temporal correlations of human pose, we incorporate the Multi-Scale Temporal Convolution (MS-TC) module Liu et al. (2020); Chen et al. (2021) into our final model. The architecture of this module is illustrated in Fig 1(b). It consists of three convolution branches, each starting with a $1\times 1$ convolution to reduce the channel dimension. This is followed by varying combinations of kernel sizes and dilations. The outputs from these convolution branches are concatenated to form the final temporal feature representation.
@@ 4.1.2 NTU RGB+D 120
Liu et al. (2019) is an extended version of NTU RGB+D. It includes an additional 57,367 skeleton sequences across 60 new action classes. Similar to the original dataset, two evaluation benchmarks are recommended: (1) Cross-Subject (X-Sub), where the 106 subjects are divided into training and testing groups, and (2) Cross-Setup (X-Set), where data from samples with even setup IDs are used for training, and data from samples with odd setup IDs are used for testing.
@@ 4.3 Comparison With the State-of-the-Art Methods
In this subsection, we compare our proposed method with a comprehensive list of state-of-the-art skeleton-based action recognition methods across all three datasets. We adopt a multi-stream fusion strategy and utilize four different modalities: joint, bone, joint motion, and bone motion. These modalities are explained as follows: The joint modality refers to the original skeleton coordinates, the bone modality represents the differential of spatial coordinates, and the joint motion and bone motion modalities capture the differential along the temporal dimension of the joint and bone modalities, respectively. The results from these four modalities are fused to produce the final outcome.
@@
For NW-UCLA, our proposed method achieved an accuracy of 96.1%, slightly worse than 96.6% of InfoGCN. This dataset presents particular challenges, as it contains fewer training samples, which makes it harder for our model to perform optimally. Furthermore, the dataset only includes 20 skeleton points, which restricts the effective utilization of hypergraph topology.
Table 2 shows the accuracy of our model across four different modalities and four test benchmarks from two datasets. The results clearly show that using bone information is better than using either joint or motion information alone. Additionally, the inclusion of motion information further improves performance in all scenarios. This is because skeleton information encompasses more hierarchical features and is more suitable for directed topological modeling. The fusion of multiple modalities leads to the best performance. This result further validates the superiority of our method.
@@ 4.4 Ablation Studies
In this subsection, we investigate the impact of different modules in our method on its overall performance. We use ST-GCNYan et al. (2018) as the baseline and analyze the contributions of individual modules by sequentially adding them. For the ablation study, all experiments are conducted on the X-Sub benchmark of NTU RGB+D 120. Only the joint modality is used as input for these experiments.
@@ 4.4.1 Hypergraph Topology Studies
For the hypergraph topology we designed, we investigated whether different hypergraph topologies might yield varying accuracy in identifying different actions. We conduct experiments on three distinct hypergraph topologies and their pairwise combinations. In addition, we compare the model parameter size and recognition accuracy across different configurations. These configurations include the fusion of data from different modalities. The experimental results are presented in Table 3.
As shown in the table, the model that integrates all three hypergraph topologies achieves the best recognition performance. Among the individual topologies, h3 demonstrates the highest accuracy when all four modalities are fused. For the two-topology combinations, the pairing of h1 and h3 achieves the best accuracy. Nonetheless, the highest recognition accuracy is consistently obtained when all three hypergraph topologies are combined.
@@ 4.4.2 Hypergraph Module Studies
To verify the effectiveness of the hypergraph module and the contribution of different components in our method, we compare our hypergraph module with configurations that do not include the hypergraph component. And we analyze the size and accuracy of our model with and without the hypergraph topology. The comparison includes four configurations: the Graph convolution module with hypergraph (HGCM) and without hypergraph(GCM), as well as the Attention module with hypergraph (HAM) and without hypergraph (AM).
As shown in the Table 4, the recognition performance of model improves progressively with the addition of components. The results demonstrate that incorporating the hypergraph module in our method significantly enhances the baseline recognition performance of model. Specifically, the inclusion of our proposed HGCM and HAM modules improves accuracy by approximately 1.9%. Although the hypergraph structure increases the parameter count of model, the overall model size remains compact, and the accuracy gains are substantial.
Notably, our HAM and HGCM capture the high-order relationships between different skeleton points. This approach highlights the advantages of using high-order relationships rather than point pairs in topological modeling. These results validate the effectiveness of our method. By leveraging hypergraph topology for high-order modeling between skeleton points, our approach leads to significant improvements in performance.
@@ 5 Conclusion
In this paper, we propose a Hypergraph Fused Graph Convolutional Network (HFGCN) for action recognition. We design different skeleton point classification strategies tailored to human motion characteristics to construct hypergraph topologies. These strategies enable the model to focus on the correlations within skeleton point groups. We further employ a Hypergraph Attention Module (HAM) to learn high-order relations between skeleton points, optimizing the topological modeling at the temporal frame level. Additionally, we introduce a Hypergraph Convolution Module (HGCM) to refine the correlations between skeleton points and skeleton point groups at the spatial channel level. Experimental results on the NTU RGB+D and NTU RGB+D 120 datasets demonstrate that the proposed HFGCN outperforms existing skeleton-based methods. However, the hypergraph in our model is manually designed based on human motion theory. Future work should focus on optimizing this aspect to achieve further improvements.